OpenStudy (anonymous):

The graph of the function f (x) = ax^2 + bx + c has a turning point at (0,2) and passes through the point (1,8). Find a,b and c. help

OpenStudy (jdoe0001):

have you covered parabolas yet?

OpenStudy (anonymous):

no i have not

OpenStudy (jdoe0001):

well, you may want to cover that first I'd think

OpenStudy (mathmale):

@malik101: By "turning point," do you by any chance mean "vertex?"

OpenStudy (mathmale):

Otherwise...what would "turning point" mean?

OpenStudy (anonymous):

The general parabola \(ax^2+bx+c\) has a vertex at \(\left(-\dfrac{b}{2a},~-\dfrac{b^2-4ac}{4a}\right)\). Given that the turning point/vertex is (0,2), you have the equations: \[\begin{cases}-\dfrac{b}{2a}=0\\\\ -\dfrac{b^2-4ac}{4a}=2\end{cases}\] You have enough info here to solve for some but not all of the coefficients. To solve for the remaining unknown, plug in the other known point, (1,8): \[\begin{align*}f(1)&=a(1)^2+b(1)+c\\8&=a+b+c\end{align*}\]